OpenStudy (anonymous):

Find the slope of the line that passes through the points (3, 6) and (5, 3).

OpenStudy (zpupster):

m=the slope= y2-y1/x2-x1 so ordered pair of (x1,y1) and (x2,y2) or (3,6) and (5,3) just plug in your values m= 3-6/5-3 m=??